Income Statement
Sales 900
COGS 550
Gross Margin 350
S&G 200
EBIT 150
Interest Exp 20
Taxes 15
Net Income 115
# Shares 400

Balance Sheet

Cash 400
Account Receivable 200
Inventory 100
Total Current Assets

700

Long Term Investments 500
Property Plant and Equipment 800
Goodwill 100
Total Assets 2100
Accounts Payable 220
Total Current Liabilities 300
Total Liabilities 1720
Stockholders' Equity
Total Stockholders' Equity 380
Net Tangible Assets 280
Cash Flow
Net Income 115
Depreciation 250
Total Cash Flow from Operating Activities 365
Capital Expenditures -250
Investments 100
Total Cash Flow from investing activities -150
Dividends Paid -200
Sale Purchase of Stock -50
Net Borrowing

200

Total Cash flows from financing activities -50
Change in case and cash equivalents 165
Solve
ROE ?
EPS ?
Current Ratio ?
Coverage Ratio ?
Debt to EBITDA ?
Cash flow ?
Free Cash Flow ?

Did the firm issue or buy back stock?

Did the firm issue or buy back debt?

Can firm cover it's dividend with FCF?
Answer #1
ROE Net income / Shareholder equity =115/380 = 30.26%
EPS Net income / Number of shares =115/400 = $0.2875
Current Ratio Current assets / Current liability =700/300 = 2.33
Coverage Ratio EBIT / Interest expense =150/20 = 7.50
Debt to EBITDA (Total liabilities - current liabilities) / (EBIT + Depreciation) =(1720-300)/(150+250) = 3.55
Cash flow Change in cash and cash equivalents 165
Free Cash Flow Cash flow from operations - capex =365-250 = 115
Q) Did the firm issue or buy back stock?
Answer : Yes firm bought back stock which is reflected in cash flow from financing activity i.e. Sale Purchase of Stock -50
Q) Did the firm issue or buy back debt?
Answer : Yes, firm issued debt which is reflected in cash flow from financing activity i.e. Net Borrowing 200
Q)Can firm cover it's dividend with FCF?
Answer : No, firm's FCF does not covers dividend as the FCF amount is 115 and dividend amount is higher than that which is 200. Hence firm can't cover it's dividend with FCF